OpenStudy (anonymous):

Is x - 8 a factor of the function f(x) = -2x3 + 17x2 - 64?

OpenStudy (anonymous):

@Nurali

OpenStudy (amistre64):

if the value that makes x-8 = 0 makes the f(x) = 0 .. then yes

OpenStudy (nurali):

Yes. When the function f(x) = –2x3 + 17x2 – 64 is divided by x – 8, the remainder is zero. Therefore, x – 8 is a factor of f(x) = –2x3 + 17x2 – 64.

OpenStudy (anonymous):

a. Yes. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is zero. Therefore, x - 8 is a factor of f(x) = -2x3 + 17x2 - 64. OR c. Yes. When the function f(x) = -2x3 + 17x2 - 64 is divided by x - 8, the remainder is not zero. Therefore, x - 8 is a factor of f(x) = -2x3 + 17x2 - 64.

OpenStudy (amistre64):

\[\frac{(x-c)~Q(x)}{(x-c)}\] \[\frac{(c-c)~Q(c)}{(c-c)}=\frac 00\]